mw104969:

CAN SOMEONE PLLZZ HELP ME!!! 5x^2+2x=0

babygirl345:

5x + 2 = 0 ⇒ x1 = −2 5 = −0.4 x = 0 ⇒ x2 = 0

mw104969:

ty

nicole122603:

thats 5x squared.. right?

mw104969:

yes

babygirl345:

yee and Np

mw104969:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @nicole122603 ok well i got x=0 and x=-2/5.... \(\color{#0cbb34}{\text{End of Quote}}\) ok but how you got a -2/5

umm:

Same as last question, starting out with factoring left side of the equation.\[x(5x+2)=0\] and then set factors that equal to 0. \[x=0 \]or\[5x+2=0\] which brings you down to \[x=0\]or\[x=\frac{ -2 }{ 5 }\]

nicole122603:

thank u for explaining it.. i stuck at explaining things..

umm:

Also, @nicole122603 - Direct answers are not allowed. Remember to walk them through the question as to why you got your answer.

nicole122603:

yeah.... i already got the warning...

mw104969:

thanks guys

umm:

Oh. Didn't know.

umm:

\(\color{#ffa366}{\text{Originally Posted by}}\) @mw104969 thanks guys \(\color{ #ffa366 }{\text{End of Quote}}\) Anytime.